OpenStudy (rebecca1588):

ANWSER NEEDED?!? What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y - 1 = -3 (x - 4) B. y + 4 = 3(x +1) C. y + 4 = -3(x + 1) D. y - 4 = -3 (x - 1)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (jingle535):

Rebbecca to figure out which one is correct put the points in the equations and see which ones come out to be true. (You know if it is true if the two sides become equal, like 0=0 or 3=3)